The distance between two stations A and B is 240km. Two trains leaved both stations at the same time and went towards each other: The A train, with 45km per hour and B train, with the speed of 35km per hour. And this is where it gets weird.. At the same time, from the station A, departed a little bird with the speed of 50km per hour, towards the B train. When the bird reached the B train, it flew back, towards the A train, reached the A train and flew back, towards the B train,etc, etc.
Question: How many kilometers the bird flew in total, until the trains finally met each other?

Answer by mananth(16949)

The distance between two stations A and B is 240km. Two trains leaved both stations at the same time and went towards each other: The A train, with 45km per hour and B train, with the speed of 35km per hour. And this is where it gets weird.. At the same time, from the station A, departed a little bird with the speed of 50km per hour, towards the B train. When the bird reached the B train, it flew back, towards the A train, reached the A train and flew back, towards the B train,etc, etc.
Question: How many kilometers the bird flew in total, until the trains finally met each other?
Looks weird but it is simple if you understand the logic
The two trains move towards each other. They meet.
time they meet = d/r
r=35+45 =80 km/h
distance =240 km
time = 240/80 = 3 hours
Time the trains take to meet is the time the bird is flying
rate = 50 km/h
d= rt
=50*3=150 km